A flywheel has a moment of inertia of 1.6 × 10−3 kg.m2
. When a constant torque is
applied, it reaches an angular velocity of 1200 rev/min in 15 s. Assuming it started from
rest, find:
i) the angular acceleration
ii) the unbalanced torque applied
iii) the angle turned through 15 s
iv) the work done on the flywheel by the torque
1
Expert's answer
2020-05-28T13:16:05-0400

As per the question,

Moment of inertia of the flywheel "(I)=1.6 \u00d7 10^{\u22123} kg.m^2"

Angular velocity "(\omega)=1200\times 2\pi rad/min =40 \pi rad/sec"

time (t)=15 sec

i) Angular acceleration "(\alpha)=\dfrac{40 \pi}{15}rad/sec^2=8 \pi/3 rad/sec^2"

ii) Torque "(\tau)=I\alpha=1.6\times 10^{-3}\times \dfrac{8 \pi}{3}=\dfrac{12.8\pi}{3}\times 10^{-3}N-m"

iii) "(\theta)=\dfrac{\omega^2}{2\alpha }=\dfrac{(40\pi)^2\times 3}{2\times 12.8\times \pi}"

"=4.68\pi rad"

iv) Work done "=\dfrac{I\omega^2}{2}=\dfrac{1.6\times 10^{-3}\times(40 \pi)^2}{2}"

"=0.64 \pi^2 J"